A Meteorite shooting vertically down into the Earth, weighing 27 pounds and traveling at 130 m/s, is found 22 cm deep in the ground. What was the magnitude of the acceleration necessary to stop the meteorite? ?
Answer #1

here,

intial speed , u = 130 m/s

s = 0.22 m

let the accelration be a

v^2 - u^2 = 2 * a * s

0 - 130^2 = 2 * a * 0.22

solving for a

a = 3.84 * 10^4 m/s^2

the magnitude of accelration is 3.84 * 10^4 m/s^2